OpenStudy (anonymous):

parametric form of r=2costheeta

OpenStudy (anonymous):

it is already a parametric form

OpenStudy (anonymous):

with parameter (\theta\)

OpenStudy (anonymous):

no it is in polar form

OpenStudy (phi):

if you stay in polar coordinates, you could write the parametric form as \[ \theta= t \\ r= 2 \cos(t) \] if you want the parametric form of this curve in rectangular coordinates, first translate from polar to rectangular coordinates, using \[ x= r \cos(\theta) \\ y= r \sin(\theta) \\ x^2 + y^2 = r^2\] multiply your equation by r: \[ r^2 = 2r \cos(\theta) \\ x^2 + y^2 = 2x\\ x^2 -2x +y^2=0\] Complete the square to get \[ (x-1)^2 + y^2 = 1 \] this is the equation of a circle with radius r=1 and center (1,0) you can write it in parametric form using \[ x= h+ r \cos(t) \\ y= k+ r \sin(t) \] where (h,k) is the center of the circle